You want to take out a $285,000 mortgage (home loan). The interest rate on the loan is 5%, and the loan is for 30 years. Your monthly payments are $1,529.94.
How much will still be owed after making payments for 15 years? $
How much will still be owed after making payments for 20 years? $
How much will still be owed after making payments for 25 years? $
Round your answers to the nearest dollar.

Answer by Theo(13342)

the actual payment is 1529.941616, rounded to 6 decimal places.
i'll assume 1529.94 as stated in the problem statement.
it's close, but not exactly right.
you can use a formula to solve this, but i'll use a calculator.
keep in mind that the calculator uses the same formula that you would use manually.
if you need to know what the formula is, let me know and i'll supply it with some instruction on how to use it.
the procedure when using the calculator is:
the interest rate per month is 5% / 12 = .41666666...%.
the number of months of the original loan is 25 * 12 = 360.
after 15 years, the number of remaining months on the loan is 360 minus 15*12 = 180.
after 20 years, the number of remaining months on the loan is 360 minus 20*12 = 120.
after 25 years, the number of remaining months on the loan is 360 minus 25*12 = 60.
using the following online calculator, find the present value of each of these scenarios.
the present value is the amount still owed on the loan after those stated years.
the results of the calculator are shown below:
the first is the present value after 360 months.

the second is the present value after 15 years.

the third is the present value after 20 years.

the fourth is the present value after 35 years.

round those present values to the nearest dollar and they should be the answer you are looking for.
i also used excel to confirm these answers are correct.
these are the results from excel.




any discrepancy between the answers from excel and the calculator can be attributed to internal rounding, either in the calculator or excel.
i can provide you with the formula to use, if you need to know that, with instructions on how to use it.
let me know is you have any questions.
theo